OpenStudy (meilendurcer):

Which point is a solution of the equation? y = –3x + 5 A. (2, 1) B. (–1, –8) C. (3, –4) D. (–2, 0)

OpenStudy (anonymous):

if you are given a list of points just plug them in and see if they work

OpenStudy (zehanz):

Just pick a random number , say 20 for x. Substitute this in the equation to get y. THe point (x, y) is a solution.

OpenStudy (abb0t):

You can also pick 19388583750 for x.

OpenStudy (zehanz):

You realize then, that there are infinitely many points that are solution of the equation. If you draw these points, you'll get a straight line...

OpenStudy (zehanz):

@abb0t: thats just as random as 20 ;)

OpenStudy (abb0t):

:P

OpenStudy (anonymous):

plug in and solve if it's equal it is a point..

OpenStudy (meilendurcer):

@ZeHanz ok i have C is that correct?

OpenStudy (zehanz):

OK, C is the point (3, -4). Set x=3: then y=-3*3+5=-9+5=-4